The bond is expected to be a 5-year, $100,000 face value, 6% bond with an effective annual yield of 5%. Interest will be payable semiannually. If all goes well, the bond will be issued on March 1, 2020, and the first interest payment date will be September 1, 2020. The bonds are expected to be callable at 102 at any time on or after March 1, 2022.

Amortize the bond using the effective-interest method.

Include an amortization schedule using the effective interest method as an exhibit to follow your memo (you do not need to consider prorations).
Answer #1
Table values are based on:
n= 10
i= 3.0%
Cash Flow Amount Present Value
Interest $100,000*3% =$3,000 $3,000*PVAF(2.50%,10) =$3,000*8.75206 =$26,256
Principal $100,000 $100,000*PVIF(2.50%,10) =$1,00,000*0.78120 =$78,120
Price of Bonds $104,376
Premium on bonds =$104,376 - $100,000 =$4,376
Amortization table
Date Interest Payment Interest expenses Premium amorrtization Bond carrying amount
01-Mar-20                      1,04,376
01-Sep-20                                                  3,000                                          2,609                                 391                      1,03,985
01-Mar-21                                                  3,000                                          2,600                                 400                      1,03,585
01-Sep-21                                                  3,000                                          2,590                                 410                      1,03,175
01-Mar-22                                                  3,000                                          2,579                                 421                      1,02,754
01-Sep-22                                                  3,000                                          2,569                                 431                      1,02,323
01-Mar-23                                                  3,000                                          2,558                                 442                      1,01,881
01-Sep-23                                                  3,000                                          2,547                                 453                      1,01,428
01-Mar-24                                                  3,000                                          2,536                                 464                      1,00,964
01-Sep-24                                                  3,000                                          2,524                                 476                      1,00,488
01-Mar-25                                                  3,000                                          2,512                                 488                      1,00,000